OpenStudy (anonymous):

The polar of any point P with respect to the circle x^2+y^2 = a^2 touches the parabola y^2=4ax. Show that P lies on the curve y^2+ax=0

OpenStudy (anonymous):

when can some help me with my geometry question pls

OpenStudy (anonymous):

finished****

OpenStudy (keygrover):

What geometry problem do you need help on chris

OpenStudy (freckles):

hey @tough3z is this calculus?

OpenStudy (anonymous):

pole and polar of parabola , 2d geometry

OpenStudy (freckles):

what can you tell me about the meaning of polar of any point P w.r.t to the circle

OpenStudy (anonymous):

polar of a point P w.r.t a conic (or a circle) is the locus of intersection of tangents ,drawn through the chord that passing through the conic and the point P

OpenStudy (freckles):

Assume a>0 x^2+y^2=a^2 has radius a y^2=4ax is a parabola open to the right the other parabola y^2=-ax is open to the left  so this is sorta what are drawing looks like